OpenStudy (anonymous):

x + y – 3z = –8 2x + 2y + z = 12 3x + y – z = –2

OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. x + y – 3z = –8 2x + 2y + z = 12 3x + y – z = –2